# Learning in complex action spaces without policy gradients

### v1

@@ 3.3 Representation learning via action-in architectures
Action-value methods are commonly employed with action-out architectures, including DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018). Conversely, policy gradient algorithms that involve $Q$ approximations resort to action-in architectures for tackling complex action spaces, such as DDPG (Lillicrap et al., 2016) and MPO (Abdolmaleki et al., 2018). Considering the specific requisites of the two families of methods in their standard forms, these are reasonable choices. In particular, standard action-value methods require evaluation of all possible actions in a given state in order to perform the maximization operation, thereby an action-out architecture is more efficient from a computational perspective. In contrast, policy gradient methods that rely on $Q$ approximation require evaluation of only one or a fixed number of actions in any given state (§3.1). Hence, using action-in architectures in the context of policy gradient methods is more computationally efficient in finite action spaces and one that functionally supports $Q$ evaluation in complex action spaces.
@@ 4 Incorporating the principles into action-value learning
In the same spirit as using an MC estimator in place of exact summation or integration over the action space in policy gradient methods (§3.1), the first principle that we consolidate into action-value learning is substituting exact maximization over the action space with a sampling-based approximation. Formally, we compute an approximation of $\max_{a}Q(s,a)$ via the steps below:
@@
where $m\geq 1$ is the number of action samples in state $s$ and $\Delta_{\text{search}}$ is a probability distribution over the generally state-conditional action space $\mathcal{A}_{s}$. Without any prior information, opting for a uniform $\Delta_{\text{search}}$ is ideal as it ensures equal sampling across all possible actions in a given state. This approach, with a constant $m$, allows for action-value learning at a fixed computational cost in arbitrarily complex action spaces (be it discrete, continuous, or hybrid).
@@
Then, we refine the $\operatorname*{arg\,max}$ approximation via Eq. 17 and update the buffer $\mathcal{B}(s_{t})\leftarrow a^{\text{max}}_{t}$. This approach does not achieve generalization across states, thus compromising its general efficacy in major ways. To enable a capacity for generalization, we resort to training a state-conditional parameterized distribution function with MLE (§2.4). In other words, we train a parametric $\operatorname*{arg\,max}$ predictor $f_{\bm{\mathbf{\theta}}}(.|s_{t})$ by employing the log-likelihood gradient (11) on the stored tuples $\{(s_{t},a_{t}^{\text{max}})\}$. Notably, this paradigm naturally supports training an ensemble of such predictors, for example based on different distributions. Therefore, we can rewrite Eq. 20 to explicitly incorporate an ensemble of $k$ parametric $\operatorname*{arg\,max}$ predictors as below:
@@ 5.1 Illustrative example
We examine two simplified variants of QMLE. The first one uses local sampling around the delta parameters for $\operatorname*{arg\,max}$ approximations that are used as targets for MLE training. Precisely, we only allow samples $\mathsf{A}_{m}$ drawn from $\delta_{\bm{\mathbf{\theta}}}(s)+\xi$, where $\delta_{\bm{\mathbf{\theta}}}$ denotes the delta-based $\operatorname*{arg\,max}$ predictor and $\xi$ is a zero-mean Gaussian noise with a standard deviation of 0.001 (cf. Eq. 21). This is akin to computing an MC approximation of $\nabla_{a}Q^{\pi}\big{(}s_{t},a\!=\!\pi(s_{t})\big{)}$ in DPG (7c). The second variant incorporates global sampling alongside local sampling, by additionally sampling from the uniform distribution of Eq. 21.
@@
Figure 3: Comparison of QMLE with learning curves of DMPO, and evaluation performances of A3C, DDPG, and D4PG after training for 100M environment steps.
@@ 5.2 Benchmarking results
With the exception of the Finger Turn Hard task, QMLE consistently performs between DDPG and D4PG. Notably, it matches or outperforms DDPG on 14 out of 16 tasks, with DDPG being the closest counterpart from the policy gradient paradigm to QMLE. Moreover, QMLE substantially exceeds the performance of A3C across all tasks. This is despite QMLE being trained on 10 to 100$\times$ fewer steps compared to A3C, DDPG, and D4PG. While QMLE competes well with DMPO in low-dimensional action spaces, it trails in higher-dimensional ones. Nonetheless, the strong performance of QMLE in continuous control tasks with up to 38 action dimensions, all without policy gradients, in and of itself testifies to the core nature of our identified principles and their adaptability to action-value methods.
@@ 6 Conclusion
In this paper, we distilled the success of policy gradient methods in complex action spaces into three core principles: MC approximation of sums or integrals, amortized maximization using a special form of MLE, and action-in architectures for representation learning and generalization over actions. We then argued that these principles are not exclusive to the policy gradient paradigm and can be adapted to action-value methods. In turn, we presented a framework for incorporating adaptations of these principles into action-value methods. To examine our arguments, we instantiated QMLE by implementing our adapted principles into approximate Q-learning with deep neural networks. Our results showed that QMLE performs strongly in continuous control problems with up to 38 action dimensions, largely outperforming its closest policy gradient counterpart DDPG. These results provided empirical support for the core nature of our identified principles and demonstrated that action-value methods could adopt them to achieve similar qualities, all without policy gradients. In a comparative study using DPG and two simplified QMLE variants, we highlighted a key limitation of policy gradients and showed how QMLE could overcome it. This study serves as a motivator for a shift from policy gradients toward action-value methods with our adapted principles. It also offers a potential explanation for the improvements observed over DDPG in our benchmarking experiments.
@@ Appendix A Q-learning with maximum likelihood estimation
To effectively manage training inference costs in QMLE, we recommend allocating a larger budget to $m_{\text{greedy}}$ than to $m_{\text{target}}$. Since $m_{\text{greedy}}$ is used at most once per interaction step, increasing it incurs relatively little computational burden. In addition, more accurate $\operatorname*{arg\,max}$ approximations during training interactions can lead to higher quality data for learning, making this increase particularly beneficial. In contrast, each training update requires $m_{\text{target}}\times N_{b}$ inferences on the target $Q$-network, where $N_{b}$ is the batch size. This makes increasing $m_{\text{target}}$ much more costly in terms of training inference costs. On that account, choosing a moderate $m_{\text{target}}$ allows for computational tractability with larger batch sizes. Remarkably, a moderate $m_{\text{target}}$ could also help reduce the overestimation of action-values (Hasselt, 2010; van Hasselt et al., 2016). Also, assigning a smaller $m_{\text{target}}$ relative to $m_{\text{greedy}}$ is further justified because target maximization benefits from additional amortization. Specifically, each time a transition is sampled from the memory buffer for experience replay, we use the previously stored $\operatorname*{arg\,max}$ approximation as a prior. This approximation is then recalibrated and updated in the memory buffer for the next time that the transition is sampled for replay.
@@ B.3 Implementation
Our QMLE implementation is based on the open-source DQN codebase by Huang et al. (2022). To support reproducibility, we will make our code publicly available upon publication.
@@ B.4 Seeds and performance
All curves report the mean undiscounted return over seeds with one standard error. Performance levels of DDPG, D4PG, and A3C represent the mean over 100 episodes per seed, after training for 100M environment steps. Table 2 details the number of seeds used for each agent in our experiments, grouped by the source of the results.
@@
QMLE 5 or 10 (depending on the task)
@@
Appendix C Supplementary results
@@
  • Figure 5 presents a comparison between QMLE and policy gradient methods that rely on action-value approximation: MPO (Abdolmaleki et al., 2018), SAC (Haarnoja et al., 2018), and TD3 (Fujimoto et al., 2018).
@@
Figure 7 shows a comparison of QMLE with QT-Opt (Kalashnikov et al., 2018) and both the discrete and continuous action variants of AQL (de Wiele et al., 2020). The baseline results are taken from de Wiele et al. (2020).
@@
Appendix D Future work
D.1 Combining with other improvements
@@
In this paper, we integrated our framework into the deep Q-learning algorithm of Mnih et al. (2015), in a proof-of-concept agent that we termed QMLE (Algorithm 1). In our benchmarking experiments, we further combined QMLE with prioritized experience replay (Schaul et al., 2016; see details in Section B). While this setup is relatively basic compared to the advancements in deep Q-learning, it served our purpose of demonstrating the general competency of action-value methods in complex action spaces without involving policy gradients. We anticipate that a purposeful integration with advancements in deep Q-learning could significantly improve the performance of our QMLE agent. For instance, fundamental methods that can be trivially combined with QMLE include $N$-step returns and distributional learning, similarly to the critics in DMPO and D4PG. Certain methods, including double Q-learning (Hasselt, 2010; van Hasselt et al., 2016) and dueling networks (Wang et al., 2016) may not be directly applicable or relevant to QMLE, underscoring the importance of careful integration. We are particularly excited about using a cross-entropy classification loss in place of regression for training $Q$ approximators (Farebrother et al., 2024), as well as combining with ideas introduced by Li et al. (2023); Schwarzer et al. (2023). Moreover, formal explorations into the space of value mappings (van Seijen et al., 2019; Fatemi & Tavakoli, 2022), particularly those that benefit $Q$-function approximation with action-in architectures, offer an intriguing direction for future work.
@@
D.2 Multiagent reinforcement learning via CTDE
Figure 8: Comparison of QMLE with linear value decomposition in a continuous variant of the “climbing” game with two agents (Claus & Boutilier, 1998). Linear value decomposition leads to a suboptimal reward of 5 due its limited representational capacity ($Q\doteq U1+U2$), whereas QMLE, by modeling the joint $Q$-function without such constraints, enables decentralized $\operatorname*{arg\,max}$ predictors that guide agents to the globally optimal reward of 11.
@@ D.2 Multiagent reinforcement learning via CTDE
D.3 Curriculum shaping through growing action spaces

### v2

@@ 3.3 Representation learning via action-in architectures
Action-value methods typically use action-out architectures—such as DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018)—while policy gradient algorithms with $Q$ approximations rely on action-in architectures to handle complex action spaces—such as DDPG (Lillicrap et al., 2016) and MPO (Abdolmaleki et al., 2018). These choices reflect the distinct requirements of each family. In particular, standard action-value methods must evaluate all possible actions per state to perform maximization, making action-out architectures the more efficient choice from a computational perspective. In contrast, policy gradient methods that rely on $Q$ approximation evaluate only one or a small set of actions in any given state (§3.1). Hence, using action-in architectures in the context of policy gradient methods is more computationally efficient in finite action spaces and one that functionally supports $Q$ evaluation in complex action spaces.
@@ 4 Incorporating the principles into action-value learning
In the same spirit as using an MC estimator in place of exact summation or integration over the action space in policy gradient methods (§3.1), the first principle that we consolidate into action-value learning substitutes exact maximization over the action space with a sampling-based approximation. Formally, we compute an approximation of $\max_{a}Q(s,a)$ via the steps below:
@@
where $m\geq 1$ is the number of action samples in state $s$ and $\Delta_{\text{search}}$ is a probability distribution over the generally state-conditional action space $\mathcal{A}_{s}$. Without any prior information, opting for a uniform $\Delta_{\text{search}}$ is ideal as it ensures equal sampling across all possible actions in a given state. This approach, with a constant $m$, allows for action-value learning at a fixed computational cost in arbitrarily complex action spaces. Sampled actions are only used internally to probe the $Q$-function for $\operatorname*{arg\,max}$ approximation at a given state, and not executed in the environment.
@@
Then, we refine the $\operatorname*{arg\,max}$ approximation via Eq. 17 and update the buffer $\mathcal{B}(s_{t})\leftarrow a^{\text{max}}_{t}$. This approach does not achieve generalization across states, thus compromising its general efficacy in major ways. To enable a capacity for generalization, we resort to training a state-conditional parameterized distribution function with MLE (§2.4). In other words, we train a parametric $\operatorname*{arg\,max}$ predictor $f_{\bm{\mathbf{\theta}}}(.|s_{t})$ by employing the log-likelihood gradient (11) on the stored tuples $\{(s_{t},a_{t}^{\text{max}})\}$. Notably, this paradigm naturally supports training an ensemble of such predictors, for example based on different distribution families. Therefore, we can rewrite Eq. 20 to explicitly incorporate an ensemble of $k$ parametric $\operatorname*{arg\,max}$ predictors as below:
@@ 5.1 Illustrative example
In this study, we present a first illustrative example of how the combination of the three core principles enables QMLE to extend action-value learning to continuous-action problems. Specifically, Principle 1 allows approximate $\operatorname*{arg\,max}$ computation over non-convex $Q$-surfaces defined on continuous (i.e. non-finite) action spaces, as demonstrated in a 2D action setting. Principle 2 facilitates iterative caching and refinement of the $\operatorname*{arg\,max}$ approximation, allowing efficient reuse during and after training. Principle 3 enables function approximation of the $Q$-function over continuous actions, which is not possible with action-out architectures such as that used in DQN. Additionally, this example highlights that, with respect to Principles 1 and 2, action-value learning can offer more flexible $\operatorname*{arg\,max}$ computation than deterministic policy gradients (DPG). As a result, QMLE is able to both mimic and surpass the behavior of DPG, depending on the choice of the action sampling scheme.
@@
We examine two simplified variants of QMLE. The first one samples around the delta parameters for $\operatorname*{arg\,max}$ approximations that are used as targets for MLE training. Precisely, we only allow samples $\mathsf{A}_{m}$ drawn from $\delta_{\bm{\mathbf{\theta}}}(s)+\xi$, where $\delta_{\bm{\mathbf{\theta}}}$ denotes the delta-based $\operatorname*{arg\,max}$ predictor and $\xi$ is a zero-mean Gaussian noise with a standard deviation of 0.001 (cf. Eq. 21). This is akin to computing an MC approximation of $\nabla_{a}Q^{\pi}\big{(}s_{t},a\!=\!\pi(s_{t})\big{)}$ in DPG (7c). The second variant additionally incorporates sampling uniformly over the full action space, corresponding to $\mathsf{A}_{m_{0}}$ samples in Eq. 21. For brevity, we refer to the first variant as using local sampling, and the second as using global sampling.
@@
Figure 3: Comparison of QMLE against its closest policy gradient counterpart, off-policy DDPG, with the canonical on-policy A3C included for reference. Performance of state-of-the-art methods DMPO and D4PG is also shown, representing a potential next frontier for QMLE. DMPO and D4PG incorporate advances such as distributional and NN-step learning, which are not used in our current QMLE implementation. Performance curves (QMLE, DMPO) show mean undiscounted return $\pm$ 1 standard error over seeds. Performance levels (A3C, DDPG, D4PG) report the mean undiscounted return, averaged over 100 evaluation episodes per seed and across seeds. The number of seeds per agent is detailed in Table 2.
@@ 5.2 Benchmarking results
[…]
Appendix C provides supplementary benchmarking results against alternative policy gradient and action-value methods.
5.3 Ablation studies
We conducted ablation experiments to assess the contribution of each principle from Section 4 to QMLE’s performance and scalability. Here, we provide a concise summary of our findings and refer the reader to Appendix D for the learning curves and additional details.
In Appendix D.1, we study the impact of the number of action samples used for approximating the value-maximizing action. We find that reducing the number of samples from 1000 to as few as 2 has negligible effect on final performance, highlighting a low sensitivity to sampling budgets. Although using significantly more samples improves accuracy, the benefits often taper off due to amortized maximization, which enables reuse of previous approximations. Consequently, one can significantly reduce inference costs by lowering the sample count without sacrificing performance, a practically valuable property for large-scale tasks or when using large models where each forward pass is computationally expensive.
In Appendix D.2, we ablate amortized maximization by removing parametric $\operatorname*{arg\,max}$ predictors and relying solely on uniform sampling. Performance degrades significantly under this condition, particularly in higher-dimensional action spaces. This study confirms the critical role of amortization when sampling budgets are limited relative to the action-space complexity, an inevitable problem scenario across real-world domains where action-space complexity quickly outpaces feasible sampling budgets.
Finally, in Appendix D.3, we replace our action-in architecture with an action-out variant. This ablation proves computationally infeasible in large discrete action spaces and underperforms even in moderate action spaces, underscoring the importance of action-in architectures for scalability and generalization.
Taken together, these ablations support that each principle individually contributes to improved scalability or performance of action-value learning.
@@ 6 Conclusion
In this paper, we distilled the success of policy gradient methods in complex action spaces into three core principles: MC approximation of sums or integrals, amortized maximization using a special form of MLE, and action-in architectures for representation learning and generalization over actions. We then argued that these principles are not exclusive to the policy gradient paradigm and can be adapted to action-value methods. In turn, we presented a framework for incorporating adaptations of these principles into action-value methods. To examine our arguments, we instantiated QMLE by implementing our adapted principles into approximate Q-learning with deep neural networks. Our results showed that QMLE performs strongly in continuous control problems with up to 38 action dimensions, largely outperforming its closest policy gradient counterpart DDPG. These results provided empirical support for the core nature of our identified principles and demonstrated that action-value methods could adopt them to achieve similar qualities, all without policy gradients. In a comparative study using DPG and two simplified QMLE variants, we highlighted a key limitation of policy gradients; namely, their reliance on local updates in continuous parameter spaces, which can lead to convergence to suboptimal solutions in multimodal or non-convex value landscapes. By contrast, QMLE leverages flexible action sampling and explicit maximization to overcome this limitation. This study serves as a motivator for a shift from policy gradients toward action-value methods with our adapted principles. It also offers a potential explanation for the improvements observed over DDPG in our benchmarking experiments.
@@ Appendix A Q-learning with maximum likelihood estimation
To effectively manage training inference costs in QMLE, we recommend allocating a larger budget to $m_{\text{greedy}}$ than to $m_{\text{target}}$. Since $m_{\text{greedy}}$ is used at most once per interaction step, increasing it incurs relatively little computational burden. In addition, more accurate $\operatorname*{arg\,max}$ approximations during training interactions can lead to higher quality data for learning, making this increase particularly beneficial. In contrast, each training update requires $m_{\text{target}}\times N_{b}$ inferences on the target $Q$-network, where $N_{b}$ is the batch size. This makes increasing $m_{\text{target}}$ much more costly in terms of training inference costs. On that account, choosing a moderate $m_{\text{target}}$ allows for computational tractability with larger batch sizes. Remarkably, a moderate $m_{\text{target}}$ could also help reduce the overestimation of action-values (van Hasselt, 2010; van Hasselt et al., 2016). Also, assigning a smaller $m_{\text{target}}$ relative to $m_{\text{greedy}}$ is further justified because target maximization benefits from additional amortization. Specifically, each time a transition is sampled from the memory buffer for experience replay, we use the previously stored $\operatorname*{arg\,max}$ approximation as a prior. This approximation is then recalibrated and updated in the memory buffer for the next time that the transition is sampled for replay.
@@ B.3 Implementation
Our QMLE implementation is based on a DQN script from CleanRL (Huang et al., 2022) and incorporates prioritized experience replay adapted from Stable Baselines (Hill et al., 2018), both available under the permissive MIT license.
To support reproducibility, we release the implementation used in our benchmarking experiments at:
https://github.com/atavakol/qmle
@@ B.4 Seeds and performance
All curves show the mean undiscounted return across seeds, with one standard error. Performance levels of DDPG, D4PG, and A3C are averaged over 100 episodes per seed, after 100M environment steps of training. Table 2 lists the number of seeds used per agent, grouped by result source.
@@
QMLE 5   - Dog and Humanoid tasks
10 - all other tasks
@@
Appendix C Supplementary benchmarking results
@@
  • Figure 5 compares QMLE with policy gradient methods that use action-value approximation: MPO (Abdolmaleki et al., 2018), SAC (Haarnoja et al., 2018), and TD3 (Fujimoto et al., 2018).
@@
Figure 7 compares QMLE with QT-Opt (Kalashnikov et al., 2018) and both the discrete and continuous action variants of AQL (de Wiele et al., 2020). The baseline results are taken from de Wiele et al. (2020).
@@
Appendix D Ablation studies
In this section, we present ablation studies to evaluate the impact of the principles in our framework on the performance of QMLE.
D.1 Approximate maximization
Figure 8 shows the learning curves for QMLE with sampling budgets of 2 and 1000. Expectedly, increasing the number of samples for $Q$-maximization improves performance by yielding more accurate estimates of the TD target and greedy actions. Nevertheless, amortization dampens the negative impact of undersampling by enabling reuse of past computations over time.
Figure 8: Comparison of QMLE with sampling budgets of 2 and 1000.
D.2 Amortized maximization
Figure 9 compares the performance of QMLE against its ablation without amortized maximization. In this experiment, QMLE employs a delta-based $\operatorname*{arg\,max}$ predictor, while its ablated variant relies solely on uniform sampling for $\operatorname*{arg\,max}$ approximation. We use the same sampling budgets of $m_{\textrm{target}}=m_{\textrm{greedy}}=2$ for both variants, with QMLE allocating its budgets equally between uniform sampling and the delta-based $\operatorname*{arg\,max}$ predictor ($\rho_{\textrm{uniform}}=\rho_{\textrm{delta}}=0.5$), and the ablated variant allocating them entirely to uniform sampling ($\rho_{\textrm{uniform}}=1$).
The action spaces range from 1-dimensional (leftmost) to 6-dimensional (rightmost) for the considered problems. The results demonstrate that amortized maximization significantly improves performance, particularly as the complexity of the action space increases.
Figure 9: Comparison of a continuous variant of QMLE with and without amortized maximization.
D.3 Action-in architecture
We compare the performance of QMLE with action-in and action-out architectures. Since action-out $Q$-approximators are not readily compatible with continuous action spaces, we examine both agents on the bang-off-bang (3 bins) discretized versions of the considered environments.
The QMLE variant with an action-in architecture employs an $\operatorname*{arg\,max}$ predictor based on a factored categorical distribution, with the same sampling budgets and uniform sampling ratio as in Table 1 but with $\rho_{\textrm{delta}}=0$ and $\rho_{\textrm{factored categorical}}=1$. On the other hand, exact maximization is performed for the ablated variant as a forward pass through an action-out $Q$-approximator collects all actions’ values in a given state. Therefore, using an action-out architecture in the ablated variant obviates the need for learned $\operatorname*{arg\,max}$ predictors or any approximate maximization altogether. That is to say, when inference with an action-out architecture is computationally feasible, performing exact maximization should also be feasible given that its cost is generally negligible compared to that of inference. This, in effect, reduces the ablated variant to DQN.
Figure 10 shows the learning curves for QMLE and DQN.
  • In lower-dimensional action spaces, such as Finger Spin and Walker Walk with 2 and 6 action dimensions respectively, where DQN is computationally tractable, both QMLE and DQN achieve similar final performance levels. However, QMLE performs more sample-efficiently due to the use of an action-in architecture, which enables generalization across actions.
  • In higher-dimensional action spaces, DQN becomes computationally intractable, resulting in out-of-memory errors or exceeding computational time constraints. In contrast, QMLE performs strongly in these environments, including Dog Walk with $3^{38}\approx 1.35\times 10^{18}$ discrete actions, underscoring the benefits of action-in architectures both in terms of computational scalability and generalization across enormous action spaces.
Figure 10: Comparison of QMLE with DQN where DQN represents the ablation of action-in architectures, and in turn all three principles, in QMLE. Dashed lines indicate out-of-memory errors or excessive computational demands for DQN.
Appendix E Future work
E.1 Combining with other improvements
@@
In this paper, we integrated our framework into the deep Q-learning algorithm of Mnih et al. (2015), in a proof-of-concept agent that we termed QMLE (Algorithm 1). In our benchmarking experiments, we further combined QMLE with prioritized experience replay (Schaul et al., 2016; see details in Section B). While this setup is relatively basic compared to the advancements in deep Q-learning, it served our purpose of demonstrating the general competency of action-value methods in complex action spaces without involving policy gradients. We anticipate that a purposeful integration with advancements in deep Q-learning could significantly improve the performance of our QMLE agent. For instance, fundamental methods that can be trivially combined with QMLE include $N$-step returns (Sutton & Barto, 2018; Hessel et al., 2018) and distributional learning (Bellemare et al., 2017), similarly to the critics in DMPO and D4PG. Certain methods, including double Q-learning (van Hasselt, 2010; van Hasselt et al., 2016) and dueling networks (Wang et al., 2016) may not be directly applicable or relevant to QMLE, underscoring the importance of careful integration. We are particularly excited about using a cross-entropy classification loss in place of regression for training $Q$ approximators (Farebrother et al., 2024), as well as combining with ideas introduced by Li et al. (2023); Schwarzer et al. (2023). Moreover, formal explorations into the space of value mappings (van Seijen et al., 2019; Fatemi & Tavakoli, 2022), particularly those that benefit $Q$-function approximation with action-in architectures, offer an intriguing direction for future work.
@@
E.2 Multiagent reinforcement learning via CTDE
@@
Figure 11: Comparison of QMLE with linear value decomposition in a continuous variant of the “climbing” game with two agents (Claus & Boutilier, 1998). Linear value decomposition leads to a suboptimal reward of 5 due to its limited representational capacity ($Q\doteq U1+U2$), whereas QMLE, by modeling the joint $Q$-function without such constraints, enables decentralized $\operatorname*{arg\,max}$ predictors that guide the agents to the globally optimal reward of 11.
E.3 Curriculum shaping through growing action spaces